An airplane trip involves four legs, with three stopover the first leg is due north 620 Km, the second leg is south east 45o for 440 Km, third leg is at 53o south of west for 550 Km, and fourth leg is due               north 20km, what is the plane total displacement and direction?      


1
Expert's answer
2021-11-04T10:16:56-0400

Determine the components of total displacement:


"d_N=620-440\sin45°-550\sin53°=-130\text{ km},\\
d_W=0-440\cos45°+550\cos53°=-111\text{ km},\\
d=\sqrt{d_N^2+d_W^2}=171\text{ km}.\\\space\\
\theta=\arctan\bigg|\frac{d_N}{d_W}\bigg|=49.5°\text{S of W. }"